# Discrete Dictionary-based Decomposition Layer for Structured Representation Learning

### v1

@@ Abstract
Neuro-symbolic neural networks have been extensively studied to integrate symbolic operations with neural networks, thereby improving systematic generalization. Specifically, Tensor Product Representation (TPR) framework enables neural networks to perform differentiable symbolic operations by encoding the symbolic structure of data within vector spaces. However, TPR-based neural networks often struggle to decompose unseen data into structured TPR representations, undermining their symbolic operations. To address this decomposition problem, we propose a Discrete Dictionary-based Decomposition (D3) layer designed to enhance the decomposition capabilities of TPR-based models. D3 employs discrete, learnable key-value dictionaries trained to capture symbolic features essential for decomposition operations. It leverages the prior knowledge acquired during training to generate structured TPR representations by mapping input data to pre-learned symbolic features within these dictionaries. D3 is a straightforward drop-in layer that can be seamlessly integrated into any TPR-based model without modifications. Our experimental results demonstrate that D3 significantly improves the systematic generalization of various TPR-based models while requiring fewer additional parameters. Notably, D3 outperforms baseline models on the synthetic task that demands the systematic decomposition of unseen combinatorial data.
@@ 1 Introduction
Compositional generalization, aiming at understanding unseen data by combining known concepts, is essential for neural networks to handle complex tasks [2, 13, 12, 15, 8, 6]. Tensor Product Representation (TPR) framework [32] facilitates this by embedding the symbolic structure of data within vector spaces, providing neural networks with compositional capabilities. Within this framework, objects are represented by a tensor product of symbolic components called roles and fillers, and these representations subsequently are superimposed to represent multiple objects in a single representation. During decoding, a specific filler is retrieved from the superimposed representation through matrix multiplication using an unbinding operator correlated to a particular role. This retrieved filler is then utilized in downstream tasks. Based on this property, TPR-based neural networks have demonstrated significant generalization and applicability in fields such as associative reasoning [27, 29], mathematical problem-solving [28], and natural language processing [9, 31, 20, 33].
@@
In this work, we propose a Discrete Dictionary-based Decomposition (D3) layer for structured representation learning within the TPR framework. D3 employs the discrete representations techniques to utilize prior knowledge acquired during training for decomposition operations. Inspired by prior discrete key-value architectures [14, 37], D3 consists of multiple dictionaries, each comprising discrete, learnable key-value pairs. Unlike prior work, each dictionary of D3 is linked explicitly to individual TPR components, such as role, filler, and unbinding operator. This design allows each dictionary to capture and store the symbolic features of its corresponding TPR components during training. D3 acts as a drop-in layer that maps input data into pre-learned symbolic features for the decomposition of TPR components through a three-step process, as illustrated in Fig. 1. First, it generates multiple queries from the input data, with each query utilized for different TPR components. Next, it identifies the nearest codebook keys within each dictionary based on these queries. Finally, D3 generates structured TPR representations by aggregating the codebook values corresponding to these keys. Moreover, D3 can be seamlessly integrated into any TPR-based model by replacing the TPR component generation layer without requiring further modifications.
@@
  • We propose a novel D3 layer to tackle the decomposition problem inherent in the TPR-based approaches. D3 leverages discrete, learnable dictionaries to enhance the decomposition capabilities of TPR-based models. By mapping input data to pre-learned symbolic features stored within the dictionaries, D3 effectively generates structured TPR representations.
@@ 2 Related Work
Research on memory networks has focused on enhancing neural network capacity by integrating external memory [35, 4, 5, 23, 26, 40]. Memory-augmented neural networks store variable lengths of sequential data in this external memory and retrieve necessary information using various addressing methods [35, 5]. These writing and reading mechanisms share many similarities with our D3 approach. However, while memory networks store input features sequentially in their memory states as a continuous stream, D3 updates symbolic feature information through gradient descent into codebook parameters within dictionaries. This distinctive characteristic allows D3 to leverage the learned symbolic features to decompose unseen data after training. In another work, Lample et al. [14] introduces a learnable key-value memory layer to improve the efficiency of the Transformer by replacing the feed-forward layer. Unlike their memory layer, D3 employs key-value pairs in dictionaries explicitly linked to individual TPR components, making it well-suited for the TPR framework.
@@ 4.1 Task
This task evaluates systematic generalization in memorizing and recalling combinatorial data [22]. It consists of a discovery phase and an inference phase. During the discovery phase, the model receives the combinatorial sequential items, each combining two symbols, $x\in X$ and $y\in Y$ where $X=X_{1}\cup X_{2}$ and $Y=Y_{1}\cup Y_{2}$. The model is then required to predict an associated $y$ when a specific $x$ is presented. The SAR task uses different combination settings between training and evaluation to target systematic generalization specifically. During training, the model learns the following combination settings: (1) $X_{1}$ and $Y_{1}$ and (2) $X_{2}$ and $Y_{2}$. At the evaluation, on the other hand, the model should generalize unseen combination settings, specifically $X_{1}$ and $Y_{2}$. In the SAR task, the TPR framework regards $x$ as the role and the unbinding operator, and $y$ as the filler. Therefore, TPR-based models should systematically decompose the combinatorial data into structured representations by mapping $x$ to the role and $y$ to the filler during the discovery phase, and mapping $x$ to the unbinding operator during the inference phase to solve this task.
@@
This task [18] is a large-scale language modeling dataset consisting of lengthy corpora from Wikipedia. Although the WikiText-103 task does not directly measure the systematic generalization of the models, it is used to evaluate the effectiveness and applicability of D3 on a large-scale task beyond relatively simple tasks.
@@ 4.3.1 Qualitative Analysis
Fig. 4(c) shows the cosine similarity between the roles during the discovery phase, and Fig. 4(c) shows the cosine similarity between the roles during the discovery phase and the unbinding operator during the inference phase. Both results demonstrate that the generated representations by D3 satisfy the TPR conditions, resulting in an accuracy of nearly 100%. We also conduct the same analysis for intermediate features, particularly query and code. Figs. 4 and 4 show that each intermediate representation complements the others to satisfy the TPR condition, indicating the effectiveness of D3.
@@
Figure 6: The mean accuracy on the SAR task for 10 seeds in the ablation study, with error bar indicating SD. The default setting uses $D_{\text{code}}$ of 64, $N_{\text{code}}$ of 64, and top-$k$ of 8. Each figure shows the experimental results for the following settings: (a) Varying $D_{\text{code}}$. (b) Varying top-$k$ with $N_{\text{code}}$ constant. (c) Varying $N_{\text{code}}$ with top-$k$ constant.
@@ Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(No. 2021R1A2C3011169) and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(No. 2022R1A5A7026673).
@@ A.4 WikiText-103 task
The WikiText-103 task [18] is a large-scale language modeling dataset consisting of lengthy corpora from Wikipedia. Although the WikiText-103 task does not directly measure the systematic generalization of the models, it is used to evaluate the effectiveness and applicability of D3 on a large-scale task beyond relatively simple tasks.
@@ Appendix C Additional Experiments
Appendix D Additional Qualitative Analysis

### v2

@@ Abstract
Neuro-symbolic neural networks have been extensively studied to integrate symbolic operations with neural networks, thereby improving systematic generalization. Specifically, Tensor Product Representation (TPR) framework enables neural networks to perform differentiable symbolic operations by encoding the symbolic structure of data within vector spaces. However, TPR-based neural networks often struggle to decompose unseen data into structured TPR representations, undermining their symbolic operations. To address this decomposition problem, we propose a Discrete Dictionary-based Decomposition (D3) layer designed to enhance the decomposition capabilities of TPR-based models. D3 employs discrete, learnable key-value dictionaries trained to capture symbolic features essential for decomposition operations. It leverages the prior knowledge acquired during training to generate structured TPR representations by mapping input data to pre-learned discrete features within these dictionaries. D3 is a straightforward drop-in layer that can be seamlessly integrated into any TPR-based model without modifications. Our experimental results demonstrate that D3 significantly improves the systematic generalization of various TPR-based models while requiring fewer additional parameters. Notably, D3 outperforms baseline models on the synthetic task that demands the systematic decomposition of unseen combinatorial data. [^1] The code of D3 is publicly available at https://github.com/taewonpark/D3
@@ 1 Introduction
Compositional generalization, aiming at understanding unseen data by combining known concepts, is essential for neural networks to handle complex tasks [2, 13, 12, 16, 8, 6]. Tensor Product Representation (TPR) framework [33] facilitates this by embedding the symbolic structure of data within vector spaces, providing neural networks with compositional capabilities. Within this framework, individual objects are decomposed at the representation level into distinct symbolic components called role-filler pairs [^2] The roles and fillers depend on the task at hand. For example, in a tree structure, the role corresponds to a position within the tree, while the filler represents the label associated with that position [34]. In associative memory, the role is analogous to an associative key, and the filler corresponds to the associative value [28, 30].. The TPR framework encodes each object by taking a tensor product of its role vector and filler vector, represented as $T=filler\otimes role$, and then superimposes them to represent multiple objects within a single representation. During decoding, the TPR framework retrieves specific fillers—essential for solving tasks—from the superimposed representation through matrix multiplication using an unbinding operator correlated to a particular role, $filler=T\cdot unbind$. This retrieved filler is then utilized in downstream tasks. Based on this property, TPR-based neural networks have demonstrated significant generalization and applicability in fields such as associative reasoning [28, 30], mathematical problem-solving [29], and natural language processing [9, 32, 21, 34].
@@
In this work, we propose a Discrete Dictionary-based Decomposition (D3) layer for structured representation learning within the TPR framework. D3 employs the discrete representations techniques to utilize prior knowledge acquired during training for decomposition operations. Inspired by prior discrete key-value architectures [14, 38], D3 consists of multiple dictionaries, each comprising discrete, learnable key-value pairs. Unlike prior work, each dictionary of D3 is linked explicitly to individual TPR components, such as role, filler, and unbinding operator. This design allows each dictionary to capture and store the discrete features of its corresponding TPR components during training. D3 acts as a drop-in layer that maps input data into pre-learned discrete features for the decomposition of TPR components through a three-step process, as illustrated in Fig. 1. First, it generates multiple queries from the input data, with each query utilized for different TPR components. Next, it identifies the nearest codebook keys within each dictionary based on these queries. Finally, D3 generates structured TPR representations by aggregating the codebook values corresponding to these keys. Moreover, D3 can be seamlessly integrated into any TPR-based model by replacing the TPR component generation layer without requiring further modifications.
@@
  • We propose a novel D3 layer to tackle the decomposition problem inherent in the TPR-based approaches. D3 leverages discrete, learnable dictionaries to enhance the decomposition capabilities of TPR-based models. By mapping input data to pre-learned discrete features stored within the dictionaries, D3 effectively generates structured TPR representations.
@@ 2 Related Work
Research on memory networks has focused on enhancing neural network capacity by integrating external memory [36, 4, 5, 24, 27, 41]. Memory-augmented neural networks store variable lengths of sequential data in this external memory and retrieve necessary information using various addressing methods [36, 5]. These writing and reading mechanisms share many similarities with our D3 approach. However, while memory networks store input features sequentially in their memory states as a continuous stream, D3 updates symbolic feature information through gradient descent into codebook parameters within dictionaries. This distinctive characteristic allows D3 to leverage the learned discrete features to decompose unseen data after training. In another work, Lample et al. [14] introduces a learnable key-value memory layer to improve the efficiency of the Transformer by replacing the feed-forward layer. Unlike their memory layer, D3 employs key-value pairs in dictionaries explicitly linked to individual TPR components, making it well-suited for the TPR framework.
@@ 4.1 Task
This task evaluates systematic generalization in memorizing and recalling combinatorial data [23]. It consists of a discovery phase and an inference phase. During the discovery phase, the model receives the combinatorial sequential items, each combining two symbols, $x\in X$ and $y\in Y$ where $X=X_{1}\cup X_{2}\cup X_{3}$ and $Y=Y_{1}\cup Y_{2}$. The model is then required to predict an associated $y$ when a specific $x$ is presented. The SAR task uses different combination settings between training and evaluation to target systematic generalization specifically. During training, the model learns the following combination settings: (1) $X_{1}$ and $Y_{1}$, (2) $X_{2}$ and $Y_{2}$, and (3) $X_{3}$ and $Y$. At the evaluation, on the other hand, the model should generalize unseen combination settings, specifically $X_{1}$ and $Y_{2}$. Additionally, the task includes a hyper-parameter $p=\frac{|X_{3}|}{|X_{2}|+|X_{3}|}$ where $|X_{i}|$ denotes the cardinality of set $X_{i}$. By adjusting $p$, this task tests the systematic generalization of models under varying levels of exposure to different symbol combinations during training. In our study, we focus solely on the most challenging setting of the SAR task ($p=0.0$), where the subset $X_{3}$ is excluded. In the SAR task, the TPR framework regards $x$ as the role and the unbinding operator, and $y$ as the filler. Therefore, TPR-based models should systematically decompose the combinatorial data into structured representations by mapping $x$ to the role and $y$ to the filler during the discovery phase, and mapping $x$ to the unbinding operator during the inference phase to solve this task.
@@
This task [19] is a language modeling dataset consisting of lengthy corpora from Wikipedia. Although the WikiText-103 task does not directly measure the systematic generalization of the models, it is used to evaluate the effectiveness and applicability of D3 on a large-scale task beyond relatively simple tasks.
@@ 4.3.1 Qualitative Analysis
Fig. 5(c) shows the cosine similarity between the roles during the discovery phase, and Fig. 5(c) shows the cosine similarity between the roles during the discovery phase and the unbinding operator during the inference phase. Both results demonstrate that the generated representations by D3 satisfy the TPR conditions, resulting in an accuracy of nearly 100%. We also conduct the same analysis for intermediate features, particularly query and code. Figs. 5 and 5 show that each intermediate representation complements the others to satisfy the TPR condition, indicating the effectiveness of D3.
@@
Figure 6: The mean accuracy on the SAR task for 10 seeds in the ablation study, with error bar indicating SD. The default setting uses $D_{\text{code}}$ of 64, $N_{\text{code}}$ of 64, and top-$k$ of 8. Each figure shows the experimental results for the following settings: (a) Varying $D_{\text{code}}$. (b) Varying $N_{\text{code}}$ with top-$k$ constant. (c) Varying top-$k$ with $N_{\text{code}}$ constant.
@@ 5 Discussion and Limitations
Motivation.
From the perspective of systematic generalization, the decomposition operations in the TPR framework can be viewed as mapping unseen data to TPR components observed during training. Motivated by this, we design a decomposition module based on discrete representations, which maps input data to discrete, learned features facilitating systematic generalization in the decomposition operations of TPR. This design choice differentiates our contribution from AID’s competitive attention-based decomposition module. Additionally, each dictionary in D3 is explicitly linked to a specific TPR component, ensuring that each dictionary is responsible solely for generating its corresponding component. The generated components are then utilized in predefined TPR operations of the TPR-based models. This design ensures that each dictionary is trained to specialize in a specific TPR component.
Interpretability.
The TPR framework decomposes data at the representation level into distinct symbols, such as role-filler pairs for encoding and unbinding operators for decoding. This characteristic enhances the interpretability of models because the relationships between roles and unbinding operators explain which parts of the input the model focuses on to predict the output. However, this interpretability is reliable only when the generated structured representations satisfy the TPR conditions. In this context, D3 enhances the interpretability of models by providing structured representations that more effectively satisfy the TPR conditions than baseline models like FWM and AID. Figs. 10 and 10 demonstrate that the representations generated by D3 better conform to the TPR conditions than those from other baseline models, supporting our claim that D3 contributes to increased interpretability.
@@
Scalability.
The scalability of D3 is inherently linked to TPR operations of baseline models since the number of dictionaries in the D3 layer aligns with the number of TPR components required for their operations. As TPR operations require increasing components to handle large datasets, our method also requires a proportional increase in dictionaries, resulting in significant computational and memory overhead. As explored in prior work, one potential solution to mitigate this issue is distributing shared dictionaries across multiple heads or layers [14]. However, this approach requires further investigation and experimentation, which we plan to research in future work.
@@ Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2021R1A2C3011169 & No. 2022R1A5A7026673 & No. RS-2022-00166735 & No. RS-2023-00218987).
@@ A.4 WikiText-103 task
The WikiText-103 task [19] is a language modeling dataset consisting of lengthy corpora from Wikipedia. Although the WikiText-103 task does not directly measure the systematic generalization of the models, it is used to evaluate the effectiveness and applicability of D3 on a large-scale task beyond relatively simple tasks.
@@ Appendix C Additional Experiments
Appendix D Additional Comparisons
In this section, we expand our comparisons to include a broader range of state-of-the-art methods, as detailed below.
sys-bAbI task.
We compare D3 to state-of-the-art methods (DAM [22] and STM [15]) on the original bAbI task. Table 9 shows that existing memory networks struggle with the sys-bAbI task, highlighting the efficacy of D3 compared to these state-of-the-art memory networks.
Table 9: The mean word error rate [%] on additional comparison of the sys-bAbI task for 10 seeds.
Model     w/o sys diff ($\downarrow$) w/ sys diff ($\downarrow$)     Gap ($\downarrow$)
DAM     0.48 $\pm$ 0.20 5.25 $\pm$ 1.64     4.77
STM     0.49 $\pm$ 0.16 4.79 $\pm$ 1.53     3.70
TPR-RNN     0.79 $\pm$ 0.16 8.74 $\pm$ 3.74     7.95
        + AID     0.69 $\pm$ 0.08 5.61 $\pm$ 1.78     4.92
        + D3     0.65 $\pm$ 0.25 3.50 $\pm$ 2.07     2.85
FWM     0.79 $\pm$ 0.14 2.85 $\pm$ 1.61     2.06
        + AID     0.45 $\pm$ 0.16 1.21 $\pm$ 0.66     0.76
        + D3 (w/o F)     0.79 $\pm$ 0.30 2.58 $\pm$ 1.12     1.79
        + D3 (w/ F)     0.75 $\pm$ 0.17 1.96 $\pm$ 0.88     1.21
Sort-of-CLEVR task.
We compare D3 to vanilla Transformer [40] and Compositional Transformer [20], designed to enhance the systematic generalization capabilities of multi-head self-attention methods. Table 10 shows that the Linear Transformer significantly degrades systematic generalization performance compared to the vanilla Transformer and the Compositional Transformer. While D3 improves the performance of the Linear Transformer from a TPR perspective, it still shows limited performance in reasoning the relationships between multiple objects (Binary and Ternary) compared to the vanilla Transformer and Compositional Transformer.
Table 10: The mean accuracy [%] on additional comparison of the sort-of-CLEVR task for 10 seeds.
Model     $D_{\text{code}}$     Unary ($\uparrow$) Binary ($\uparrow$) Ternary ($\uparrow$)
Transformer     -     97.4 $\pm$ 3.5 84.3 $\pm$ 4.3 62.7 $\pm$ 3.9
Compositional Transformer     -     98.9 $\pm$ 0.2 88.4 $\pm$ 1.4 66.5 $\pm$ 1.9
Linear Transformer     -     69.3 $\pm$ 14.8 75.5 $\pm$ 1.3 56.4 $\pm$ 4.3
            + AID     -     98.9 $\pm$ 0.2 78.6 $\pm$ 0.3 63.7 $\pm$ 1.2
            + D3 (w/o F)     128     73.9 $\pm$ 16.5 77.2 $\pm$ 2.2 57.3 $\pm$ 4.6
    256     73.7 $\pm$ 16.5 77.8 $\pm$ 2.5 57.9 $\pm$ 5.8
            + D3 (w/ F)     128     98.9 $\pm$ 0.2 79.5 $\pm$ 0.8 63.1 $\pm$ 1.9
    256     99.0 $\pm$ 0.3 82.1 $\pm$ 2.4 68.8 $\pm$ 1.2
WikiText-103 task.
We compared D3 to the Delta Network [31], which introduced a delta updating rule instead of the additive outer product-based updating rule in the Linear Transformer. Table 11 indicates that although D3 improves the performance of the Linear Transformer in language modeling tasks, the choice of updating rules has a more substantial impact on performance for tasks involving the comprehension of lengthy corpora than the decomposition operation.
Table 11: Perplexity on additional comparison of the WikiText-103 task.
Model     $D_{\text{code}}$     Valid ($\downarrow$) Test ($\downarrow$)
Delta Network     -     35.640 36.659
Linear Transformer     -     36.473 37.533
             + AID     -     36.159 37.151
             + D3 (w/o F)     32     36.061 37.220
    64     35.975 37.009
             + D3 (w/ F)     32     36.630 37.620
    64     36.220 37.128
Appendix E Additional Ablation Study
In this section, we extend our ablation studies to investigate the effects of varying the number of keys in the codebook and the impact of removing either the residual connection or the codebook from the D3 layer.
The Effect of Varying the Number of Codebook Keys.
Fig. 7 shows that even with a significantly reduced number of keys, the model with D3 maintains high accuracy on the SAR task. This observation prompts the question of how consistent performance is achieved despite the reduction in codebook size. To explore this further, we examine the impact of removing the codebook or the residual connection within the D3 layer on the SAR and sys-bAbI tasks. Specifically, removing the codebook means that the components are generated solely by the shared feed-forward networks ($\text{layer}_{\text{residual}}$ and $\text{layer}_{\text{final}}$) while removing the residual connection implies that the components are derived solely from the codebook values.
Figure 7: The mean accuracy on the SAR task for 10 seeds in the ablation study for the effect of varying $N_{\text{code}}$ from 2 to 128 with top-$k$ constant.
The Effect of Residual Connection.
Fig. 8 shows that without the residual connection, the generalization performance of D3 dramatically degrades. This result indicates that the residual connection is crucial for effectively training the D3 layer.
Figure 8: Ablation study for the effect of the residual connection on (a) the SAR task and (b) the sys-bAbI task for 10 seeds.
The Effect of Codebook.
Table 13 shows that even without the codebook ("w/o codebook"), the D3 layer improves the generalization performance of the baseline model on the SAR task. This result indicates that the shared feed-forward networks significantly contribute to performance enhancement, which may explain why the model maintains robust performance even with fewer keys.
However, it is important to note that without the codebook, the D3 layer does not achieve near-perfect accuracy on the SAR task (as shown in Table 13) and fails to significantly enhance the systematic generalization of the baseline model on the sys-bAbI task (as shown in Table 13). These results demonstrate that the codebook plays a crucial role in enhancing the model’s overall performance and generalization capabilities, especially in tasks requiring systematic generalization.
Furthermore, we experiment with $N_{\text{code}}=1$ on the SAR task, where the codebook may act as a bias term. The results in Table 13) show that using a single codebook element leads to degraded generalization performance compared to the "w/o codebook" configuration, indicating that multiple codebook elements are essential for achieving optimal results.
Table 12: Ablation study for the effect of the codebook on the SAR task for 10 seeds.
Model     $D_{\text{code}}$ $N_{\text{code}}$ top-$k$     Accuracy ($\uparrow$)
FWM     - - -     44.90 $\pm$ 31.5
        + D3     4 2     87.38 $\pm$ 11.10
        + D3     64 8     99.27 $\pm$ 0.88
        + D3 (w/o codebook)     32 - -     89.02 $\pm$ 4.56
        + D3     1 1     89.10 $\pm$ 7.99
        + D3     4 2     94.47 $\pm$ 2.35
        + D3     64 8     94.29 $\pm$ 8.06
        + D3 (w/o codebook)     64 - -     91.65 $\pm$ 3.66
Table 13: Ablation study for the effect of the codebook on the sys-bAbI task for 10 seeds.
Model     w/o sys diff ($\downarrow$) w/ sys diff ($\downarrow$)     Gap ($\downarrow$)
FWM     0.79 $\pm$ 0.14 2.85 $\pm$ 1.61     2.06
        + D3     0.75 $\pm$ 0.17 1.96 $\pm$ 0.88     1.21
        + D3 (w/o codebook)     1.19 $\pm$ 0.41 3.55$\pm$ 1.04     2.36
Discussion.
Our ablation study on the codebook in the SAR task (Table 13) indicates that the shared residual networks within the D3 layer significantly enhance generalization performance. However, the results from the sys-bAbI task (Table 13) suggest that while these networks improve performance, they alone struggle to generalize more structured data.
The ablation studies in Tables 13 and 13 demonstrate that incorporating the codebook mechanism leads to nearly 100% accuracy on the SAR task and significantly improves the systematic generalization of models on the sys-bAbI task. However, as shown in the ablation study on the residual connection (Fig. 8), the codebook alone does not achieve the same level of generalization and exhibits instability within the D3 layer.
In conclusion, our experimental results indicate that the combination of the codebook and the shared residual networks within the D3 layer is crucial for enhancing systematic generalization performance and stability. By integrating these two components, our D3 layer significantly improves the systematic generalization capabilities of TPR-based models.
Appendix F Additional Qualitative Analysis
F.1 Comparison to Baselines
We conduct an orthogonal analysis for the baseline models (FWM and AID) similar to the analysis presented in Section 4.3.1. Figs. 10 and 10 indicate that the D3 model generates more structured and orthogonal representations than the baseline models, FWM and AID, demonstrating its effectiveness.
Figure 9: The heatmap displays the cosine similarity between the roles during the discovery phase for the SAR task.
Figure 10: The heatmap displays the cosine similarity between the roles (x-axis) during the discovery phase and the unbinding operators (y-axis) during the inference phase for the SAR task.
F.2 Qualitative Analysis for Different Seeds